The perimeter of a rectangle is 28 inches.
The length is 6 less than 3 times the
width. What is the length of the
rectangle?

Answer by TimothyLamb(4379)

2L + 2W = 28
L = 3W - 6
---
2L + 2W = 28
2(3W - 6) + 2W = 28
6W - 12 + 2W = 28
8W = 28 + 12
W = (28 + 12)/8
W = 5
L = 3W - 6
L = 3*5 - 6
L = 9
---
answer:
L = 9 inches
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php